The length of the top of a table is 5m greater than the width. The are is 84m^2. Find the dimensions of the table
Answer by waynest(281)

length = w + 5
A = lw
84 = (w + 5)w
84 = w^2 + 5w
y = w^2 + 5w -84
(w - 7)(w + 12)
w - 7 = 0
w - 7 + 7 = 7
w = 7
w + 12 = 0
w + 12 - 12 = -12
w = -12
w = 7 is correct, you can not have a negative value for the dimension.


check:
84 = (7 + 5)7
84 = (12)7
84 = 84